OpenStudy (anonymous):

the triangle has vertices D(-1,6), E(4,3), F(0,-4). draw an altitude from vertex D. Find an equation for the altitude from vertex D.

OpenStudy (anonymous):

its a right angle triangle

OpenStudy (anonymous):

so the side DE is the altitude

OpenStudy (amistre64):

y = (3x-12 +15)/5 y = \(-\frac{3x+3}{5}\) if i did it right

OpenStudy (amistre64):

and I dint.... bummer

OpenStudy (amistre64):

y = -3/5 (x-4)+3 this part is right :)

OpenStudy (anonymous):

the final answer given is y= - 4/7x+ 38/7

OpenStudy (amistre64):

y = (-3/5)x +(12/5) +3 oh; then DE isnt the altitude

OpenStudy (anonymous):

but its a right angle triangle and its given the altitude is DE

OpenStudy (amistre64):

id have to draw it out to see it correctly;

OpenStudy (anonymous):

k..i tried it though..and yes it is a right angle triangle

OpenStudy (amistre64):

none of the slopes between lines are perpendicular; and none of the sides that I find fit the pytahg; i cant see it as a 90 degree triangle

OpenStudy (amistre64):

slope are: -10; 7/4; and -5/3 in order for this to be a right tri; 2 of these whould have to multiply to be -1

OpenStudy (anonymous):

how do we get that

OpenStudy (amistre64):

to find the altitude of d; we determine the slope of ef and take its reciprocal and negate it for our slope 7/4; becomes -4/7 now input out d point like this: y = (-4/7)(x+1)+6 and solve

OpenStudy (amistre64):

i get: y = (-4/7)x - (38/7)

OpenStudy (amistre64):

err... + 38/7 on the end; confused my signs

OpenStudy (anonymous):

k..so we shld substitute the point (-1,6)??

OpenStudy (amistre64):

yes y = slope(x-Px)+Py slope = -4/7 ; Px = -1; Py = 6

OpenStudy (anonymous):

and y shld we take the reciprocal?

OpenStudy (amistre64):

becasue perpendicular lines have that property to them; their slopes multiply to -1 if we have a slope of 7/4 we need to flip it to get 1 \[\frac{7}{4}*\frac{4}{7}=1\] but we need it to be \(-1\) which is why we negate it after we flip it over

OpenStudy (amistre64):

gotta head to class; have fun and good luck :)

OpenStudy (anonymous):

thank u so much..:)